Question: Describe 2 ways you could set up a 1:10,000 dilution of a culture.

This is in regards to the Pour Plate Method of isolation and serial dilutions for microbiology. I think I may be overthinking this. Is it just using 4 tubes with 10ml each or 3 tubes, one with 100ml and two with 10ml? Transferring 1ml each time.Thank you in advance for your help.
Answer #1

First way:

Take 1ml from the solution in tube 1 and add in tube 2 and dissolve 9 ml of the water in it. In this way solution becomes 1:10

Now from tube 2, take 1 ml in tube 3 and dissolve 9 ml of water.

The dilution becomes 1/10×1/10=1/100

Now take 1 ml from tube 3 and add in tube 4 and dissolve 9 ml water. The serial dilution becomes 1/100×1/10= 1/1000.

Way 2:

Take 1 ml from 10 ml test tube 1 and add in 10 ml if test tube 2. Now dissolve 9 ml of water in it. The dilution becomes 1/10

Now take 1 ml from test tube 2 and add in 100 ml of test tube 3. Now dissolve 99 ml of water in it. In this way serial dilution becomes 1/10×1/100=1/1000.